### v2

@@
Spanning Trees Minimizing Branching Costs ***An extended abstract of some of the results of this paper appeared in [23].
[…]
Department of Computer Science
University of Salerno
Italy
{lgargano, arescigno}@unisa.it
@@
Keywords: Spanning tree, Fixed parameterized algorithms, Modular-width, Neighborhood diversity
Declarations
  • Funding: This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU - NGEU.
  • Competing interests: The authors have no competing interests to declare that are relevant to the content of this article.
@@ 1 Introduction
The problem of determining a spanning tree with a limited number of branch vertices, though fundamentally a theoretical question, originated from addressing challenges in wavelength-division multiplexing (WDM) technology within optical networks. WDM is a highly effective method for utilizing the full bandwidth potential of optical fibers, thereby meeting the high bandwidth demands of the Internet [28].
@@
Multicasting involves the simultaneous transmission of information from a single source to multiple destinations. In Wavelength-Division Multiplexing (WDM) systems, light trees are crucial for efficiently managing multicast communications by duplicating and routing specific wavelength channels. Light-splitting switches in these systems separate individual wavelengths within an optical signal. Due to the high cost of light-splitting switches, it is essential to minimize the number of nodes equipped with these devices. This need gives rise to the Minimum Branch Vertices (MBV) problem, which seeks to reduce the number of light-splitting switches in a light tree[20]. In Cognitive Radio Networks, as well as in 5G technologies that operate across a broad spectrum of frequencies, managing the switching costs associated with transitions between different service providers is of utmost importance. This is essential not only for minimizing delays but also for optimizing energy consumption [27, 37].
@@
The Minimum Branch Vertices (MBV) problem has been extensively studied from both algorithmic and graph-theoretic perspectives. Most prior research has focused on establishing upper bounds for the number of branch vertices in the resulting tree, though these bounds were often not tight. Gargano et al. [21] proved that determining whether a graph $G$ has a spanning tree with at most $k$ branch vertices is NP-complete, for each $k\geq 1$, even in cubic graphs. In the same paper, they provided an algorithm that finds a spanning tree with one branch vertex if every set of three independent vertices in $G$ has a degree sum of at least $|V(G)|-1$. Salamon [36] developed an algorithm that finds a spanning tree with $O(\log|V(G)|)$ branch vertices for graphs where each vertex has a degree of $\Omega(n)$; moreover, an approximation factor better than $O(\log|V(G)|)$ would imply that $NP\subseteq DTIME(n^{O(\log\log n)})$. Sufficient conditions for a connected claw-free graph to have a spanning tree with $k$ branch vertices are presented in [34].
@@
Integer linear formulations of the MBV problem and its variants are discussed in [9, 10, 11], including different relaxations and numerical comparisons of these relaxations. Hybrid integer linear programs for MBV are considered and solved with branch-and-cut algorithms in [39]. Decomposition methods for solving the MBV problem are explored in [31, 33, 35], while other heuristics are presented in [32, 38, 40].
@@
A complementary formulation called the Maximum Path-Node Spanning Tree (MPN), which aims to maximize the number of vertices with a degree of at most two, is investigated in [12]. The authors prove that MPN is APX-hard and provide an approximation algorithm with a ratio of 6/11. Related gathering processes are considered in [5, 6, 7, 24, 25].
@@
Notice that the Partitioning into Paths problem was already considered in [18]; however as originally defined, it only asks for the value $ham(G)$, while we ask for the actual path partitioning of $G$. Recently, the PP problem and some its variants have been considered in [16].
@@ 1.2 Our Results
Additionally, in Section 4 we direct our focus to a scenario where the cost associated with selecting a vertex as a branch vertex varies across the graph. Consequently, we delve into the case where each vertex $v\in V$ is assigned a specific non-negative integer weight $w(v)$ and the cost of a spanning tree $T$ is $w(T)=\sum w(u)$, where the sum is over all the branch vertices of $T$. The goal of the weighted version of the MBV is to find a spanning tree $T$ of the input graph that minimizes $w(T)$. For the weighted MBV problem we present an FPT algorithm parameterized by neighborhood diversity. An FPT algorithm for the case of uniform cost was given in [22].
@@
It was recently proven that MBV is FPT when parameterized by treewidth [4]. The algorithm presented in [4] runs in time $O(4^{2{\tt tw}}{\tt tw}^{4{\tt tw}+1}({\tt tw}+1)\;n)$, where tw is the treewidth of the input graph. On the other hand, it was shown in [17] that the problem is $W[1]$-hard when parameterized by clique-width. Specifically, in [17] it was proven that the (MBV special case) hamiltonian path problem is $W[1]$-hard when parameterized by clique-width.
@@
Modular-width. Modular-width was introduced in [18] and is defined as follows.
@@
We recall that an algebraic expression of width ${\tt{mw}}({{H}})$ can be constructed in linear time [41].
@@
Neighborhood diversity. The neighborhood diversity parameter, was first introduced by Lampis in [30]. It has then received much attention [1, 2, 8, 13, 14, 15, 19, 24, 26, 42], also due to the fact that, contrary to other popular parameters, it is computable in linear time.
@@ 2 The MBV problem
    Note that the aforementioned step is facilitated by (1). It’s important to emphasize that even if $w$ becomes a new branch vertex, we observe that $|V(G_{j})\cap B|=1$. Furthermore, $|V(G_{i})\cap B|\leq 1$, ensuring that the total count of branch vertices remains unchanged. Consequently, we have successfully derived a new spanning tree $T^{\prime}$ with a set of branch vertices $B^{\prime}$ such that $|B^{\prime}|\leq|B|$. By repeating steps a) and b), one can construct the desired spanning tree of ${{G}}$ with at most one branch vertex in each $V(G_{i})$. ∎
In the remaining part of this section, we present an algorithm that computes a spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with $b({{G}})$ branch vertices, if $b({{G}})>0$. In Section 3.2 we deal with the case $b({{G}})=0$, that is, we show how to find a Hamiltonian path of ${{G}}$, if one exists. The work in [18] focuses on verifying the existence of a Hamiltonian path in a graph; the proofs therein implicitly give a method for constructing such a path. We recall such a construction and show its relationship as a special case of our algorithm.
@@ 2.2 The spanning tree construction
The EXPLORE$(u)$ procedure initiates exploration from $u=f(P^{j}_{k})$ along with the entire path or spider $P^{j}_{k}$ Assume that when in the algorithm $P\in{\cal{P}}^{j}$ is explored (i.e., $Ex=Ex\cup P$), the parent function $\pi$ is updated. For path, $\pi$ is set from $f(P)$ to $s(P)$, and for spiders, it is set from the center $f(P)=s(P)$ to the leaves.. Initially, only the vertex $s(P^{j}_{k})$ is placed in $Q$. Subsequently, the main tree $T$ is expanded. If $u\neq f(P^{r}_{1})$, the parent of $u$ is already a vertex in $T$, facilitating the construction of a subtree rooted at $u$ that spans all newly explored vertices.
@@ 2.3 The algorithm complexity
Jansen and Rohwedderb [29] have recently showed that the time needed to find a feasible solution of an ILP with $p$ integer variables and $q$ constraints is $O(\sqrt{q}\Delta)^{(1+o(1))q}+O(qp)$, where $\Delta$ is the largest absolute value of any coefficient in the ILP. Denoted by ${{m}}$ the number of edges of ${{H}}$, our ILP has $q=3{{n}}+2{{m}}+1$ constraints, $p=2({{m}}+1)$ variables and $\Delta={{n}}$. Hence the time to solve it is $O({{n}}\sqrt{{{n}}+{{m}}})^{(1+o(1))(3{{n}}+2{{m}}+1)}+O({{n}}({{n}}+{{m}})).$ Using the solution $(y,x)$ of the ILP, the algorithm TREE returns the spanning tree of ${{G}}$ in time $O(|V({{G}})|^{2})$. Overall, the algorithm requires time
@@ 3 The triple and partition computation
Following [18], we use a bottom-up dynamic programming approach along the parse-tree of an algebraic expression describing ${{G}}$, to compute for every internal vertex ${\hat{{{G}}}}={\hat{{{H}}}}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}})$ of the parse tree a record of data, using those already computed for its children. Specifically, given the triple $(ham({\hat{G}}_{i}),spi({\hat{G}}_{i}),|V({\hat{G}}_{i})|)$,  ${\cal{P}}_{ham({\hat{G}}_{i})}$ and ${\cal{P}}_{spi({\hat{G}}_{i})}$, for each $i=1,\ldots,{\hat{n}}$, we need to compute
@@ 4 The weighted MBV  problem
Trivially, in case $w(u)=1$ for each $u\in V$ then CBV  becomes the problem of determining a spanning tree having the minimum number of branch vertices (MBV).
@@
The time complexity of the algorithm can be easily derived by considering that: The ILP is solved at most for each $B_{H}\subseteq\{1,\cdots,{\tt{nd}}\}$ and, that it has at most $q={\tt{nd}}^{2}+3{\tt{nd}}+2$ constraints, at most $p=2({\tt{nd}}^{2}+1)$ variables and $\Delta={\tt{nd}}$; the algorithm TREE takes time $O(V(G)^{2})$. By the results of Jansen and Rohwedderb [29], overall the algorithm requires time

### v3

@@
Abstract
@@
keywords:
Spanning tree, Fixed parameterized algorithms, Modular-width, Neighborhood diversity
@@ 1 Introduction
The problem of determining a spanning tree with a limited number of branch vertices, though fundamentally a theoretical question, originated from addressing challenges in wavelength-division multiplexing (WDM) technology within optical networks. WDM is a highly effective method for utilizing the full bandwidth potential of optical fibers, thereby meeting the high bandwidth demands of the Internet He et al. (2002).
@@
Multicasting involves the simultaneous transmission of information from a single source to multiple destinations. In Wavelength-Division Multiplexing (WDM) systems, light trees are crucial for efficiently managing multicast communications by duplicating and routing specific wavelength channels. Light-splitting switches in these systems separate individual wavelengths within an optical signal. Due to the high cost of light-splitting switches, it is essential to minimize the number of nodes equipped with these devices. This need gives rise to the Minimum Branch Vertices (MBV) problem, which seeks to reduce the number of light-splitting switches in a light tree Gargano et al. (2004). In Cognitive Radio Networks, as well as in 5G technologies that operate across a broad spectrum of frequencies, managing the switching costs associated with transitions between different service providers is of utmost importance. This is essential not only for minimizing delays but also for optimizing energy consumption Gozupek et al. (2013); Shami and Rasti (2016).
@@
The Minimum Branch Vertices (MBV) problem has been extensively studied from both algorithmic and graph-theoretic perspectives. Most prior research has focused on establishing upper bounds for the number of branch vertices in the resulting tree, though these bounds were often not tight. Gargano et al. (2002) proved that determining whether a graph $G$ has a spanning tree with at most $k$ branch vertices is NP-complete, for each $k\geq 1$, even in cubic graphs. In the same paper, they provided an algorithm that finds a spanning tree with one branch vertex if every set of three independent vertices in $G$ has a degree sum of at least $|V(G)|-1$. Salamon (2005) developed an algorithm that finds a spanning tree with $O(\log|V(G)|)$ branch vertices for graphs where each vertex has a degree of $\Omega(n)$; moreover, an approximation factor better than $O(\log|V(G)|)$ would imply that $NP\subseteq DTIME(n^{O(\log\log n)})$. Sufficient conditions for a connected claw-free graph to have a spanning tree with $k$ branch vertices are presented in Matsuda et al. (2014).
@@
Integer linear formulations of the MBV problem and its variants are discussed in Carrabs et al. (2013); Cerrone et al. (2014); Cerulli et al. (2009), including different relaxations and numerical comparisons of these relaxations. Hybrid integer linear programs for MBV are considered and solved with branch-and-cut algorithms in Silvestri et al. (2017). Decomposition methods for solving the MBV problem are explored in Landete et al. (2017); Melo et al. (2016); Rossi et al. (2014), while other heuristics are presented in Marin (2015); Silva et al. (2014); Sundar et al. (2012).
@@
A complementary formulation called the Maximum Path-Node Spanning Tree (MPN), which aims to maximize the number of vertices with a degree of at most two, is investigated in Chimani and Spoerhase (2015). The authors prove that MPN is APX-hard and provide an approximation algorithm with a ratio of 6/11. Related gathering processes are considered in Bermond et al. (2008, 2010, 2013); Gargano and Rescigno (2015, 2009).
@@
Notice that the Partitioning into Paths problem was already considered in Gajarský et al. (2013); however as originally defined, it only asks for the value $ham(G)$, while we ask for the actual path partitioning of $G$. Recently, the PP problem and some its variants have been considered in Fernau et al. (2023).
@@ 1.2 Our Results
Additionally, in Section 4 we direct our focus to a scenario where the cost associated with selecting a vertex as a branch vertex varies across the graph. Consequently, we delve into the case where each vertex $v\in V$ is assigned a specific non-negative integer weight $w(v)$ and the cost of a spanning tree $T$ is $w(T)=\sum w(u)$, where the sum is over all the branch vertices of $T$. The goal of the weighted version of the MBV is to find a spanning tree $T$ of the input graph that minimizes $w(T)$. For the weighted MBV problem we present an FPT algorithm parameterized by neighborhood diversity. An FPT algorithm for the case of uniform cost was given in Gargano and Rescigno (2023b).
@@
It was recently proven that MBV is FPT when parameterized by treewidth Baste and Watel (2022). The algorithm presented in Baste and Watel (2022) runs in time $O(4^{2{\tt tw}}{\tt tw}^{4{\tt tw}+1}({\tt tw}+1)\;n)$, where tw is the treewidth of the input graph. On the other hand, it was shown in Fomin et al. (2009) that the problem is $W[1]$-hard when parameterized by clique-width. Specifically, in Fomin et al. (2009) it was proven that the (MBV special case) hamiltonian path problem is $W[1]$-hard when parameterized by clique-width.
@@
Modular-width. Modular-width was introduced in Gajarský et al. (2013) and is defined as follows.
@@
We recall that an algebraic expression of width ${\tt{mw}}({{H}})$ can be constructed in linear time Tedder et al. (2008).
@@
Neighborhood diversity. The neighborhood diversity parameter, was first introduced by Lampis in Lampis (2012). It has then received much attention Agrawal et al. (2020); Araujo et al. (2018); Bhyravarapu and Reddy (2022); Cordasco et al. (2018, 2020, 2022); Ganian (2012); Gargano and Rescigno (2015); Gavenciak et al. (2022); Toufar et al. (2019), also due to the fact that, contrary to other popular parameters, it is computable in linear time.
@@ 2 The MBV problem
    Note that the aforementioned step is facilitated by (1). It’s important to emphasize that even if $w$ becomes a new branch vertex, we observe that $|V(G_{j})\cap B|=1$. Furthermore, $|V(G_{i})\cap B|\leq 1$, ensuring that the total count of branch vertices remains unchanged. Consequently, we have successfully derived a new spanning tree $T^{\prime}$ with a set of branch vertices $B^{\prime}$ such that $|B^{\prime}|\leq|B|$. By repeating steps a) and b), one can construct the desired spanning tree of ${{G}}$ with at most one branch vertex in each $V(G_{i})$.
In the remaining part of this section, we present an algorithm that computes a spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with $b({{G}})$ branch vertices, if $b({{G}})>0$. In Section 3.2 we deal with the case $b({{G}})=0$, that is, we show how to find a Hamiltonian path of ${{G}}$, if one exists. The work in Gajarský et al. (2013) focuses on verifying the existence of a Hamiltonian path in a graph; the proofs therein implicitly give a method for constructing such a path. We recall such a construction and show its relationship as a special case of our algorithm.
@@ 2.2 The spanning tree construction
The EXPLORE$(u)$ procedure initiates exploration from $u=f(P^{j}_{k})$ along with the entire path or spider $P^{j}_{k}$  [^1] Assume that when in the algorithm $P\in{\cal{P}}^{j}$ is explored (i.e., $Ex=Ex\cup P$), the parent function $\pi$ is updated. For path, $\pi$ is set from $f(P)$ to $s(P)$, and for spiders, it is set from the center $f(P)=s(P)$ to the leaves.. Initially, only the vertex $s(P^{j}_{k})$ is placed in $Q$. Subsequently, the main tree $T$ is expanded. If $u\neq f(P^{r}_{1})$, the parent of $u$ is already a vertex in $T$, facilitating the construction of a subtree rooted at $u$ that spans all newly explored vertices.
@@ 2.3 The algorithm complexity
Jansen and Rohwedderb (2023) have recently showed that the time needed to find a feasible solution of an ILP with $p$ integer variables and $q$ constraints is $O(\sqrt{q}\Delta)^{(1+o(1))q}+O(qp)$, where $\Delta$ is the largest absolute value of any coefficient in the ILP. Denoted by ${{m}}$ the number of edges of ${{H}}$, our ILP has $q=3{{n}}+2{{m}}+1$ constraints, $p=2({{m}}+1)$ variables and $\Delta={{n}}$. Hence the time to solve it is $O({{n}}\sqrt{{{n}}+{{m}}})^{(1+o(1))(3{{n}}+2{{m}}+1)}+O({{n}}({{n}}+{{m}})).$ Using the solution $(y,x)$ of the ILP, the algorithm TREE returns the spanning tree of ${{G}}$ in time $O(|V({{G}})|^{2})$. Overall, the algorithm requires time
@@ 3 The triple and partition computation
Following Gajarský et al. (2013), we use a bottom-up dynamic programming approach along the parse-tree of an algebraic expression describing ${{G}}$, to compute for every internal vertex ${\hat{{{G}}}}={\hat{{{H}}}}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}})$ of the parse tree a record of data, using those already computed for its children. Specifically, given the triple $(ham({\hat{G}}_{i}),spi({\hat{G}}_{i}),|V({\hat{G}}_{i})|)$,  ${\cal{P}}_{ham({\hat{G}}_{i})}$ and ${\cal{P}}_{spi({\hat{G}}_{i})}$, for each $i=1,\ldots,{\hat{n}}$, we need to compute
@@ 4 The weighted MBV  problem
. Trivially, in case $w(u)=1$ for each $u\in V$ then CBV  becomes the problem of determining a spanning tree having the minimum number of branch vertices (MBV).
@@
The time complexity of the algorithm can be easily derived by considering that: The ILP is solved at most for each $B_{H}\subseteq\{1,\cdots,{\tt{nd}}\}$ and, that it has at most $q={\tt{nd}}^{2}+3{\tt{nd}}+2$ constraints, at most $p=2({\tt{nd}}^{2}+1)$ variables and $\Delta={\tt{nd}}$; the algorithm TREE takes time $O(V(G)^{2})$. By the results of Jansen and Rohwedderb Jansen and Rohwedderb (2023), overall the algorithm requires time